OpenStudy (nikolas):

Consider the following equilibrium: | H2(g) + CO2(g) <-> CO(g) + H2O(g), Keq=0.12 | Initially, 1.0mol of CO and 1.0mol H2O are placed in a 2.0L container. Calculate the equilibrium [CO].

OpenStudy (nikolas):

I tried making an ICE chart to find [H2] and [Co2] = x while [CO] and [H2O] = 0.5-x, and then using \[\sqrt{0.12}=((.5-x)/x)\] but I keep getting that x = 0.7M when the answer is that [CO]=0.13M

OpenStudy (matt101):

The first part of your process is correct. However, by taking the square root of both sides right away, you convert a quadratic equation into a linear one. In other words, you eliminate one solution for x completely. If you solve the equation without taking the square root, you'll get two values of x: 0.76 and 0.37. x cannot be 0.76 since that is greater than the concentration you start with, so it must be 0.37. 0.5 - 0.37 = 0.13, your equilibrium concentration of CO (and H2O).

OpenStudy (nikolas):

Thank you! Not sure why I did that immediately, but all of the problems make much more sense now =)